OpenStudy (anonymous):

Solve using quadratic formula: (x+5)(x-6)=-26 simplify, use radicals as needed, express complex numbers in terms of i.

OpenStudy (anonymous):

(x+5)(x-6) =-26 x^2-x -30 = -26 x^2 -x -4 = 0 x= (1+- sqrt(1+16))/2 x = (1+ sqrt(17))/2 or (1-sqrt(17))/2

OpenStudy (saifoo.khan):

\[\Large x = \frac{1}{2}(1-\sqrt{17}),\frac{1}{2}(1+\sqrt{17})\]